A hydroelectric power station generates electricity by using 500 kg s-1 of water which falls from the storage reservoir through a height of 7.5m before flowing over the turbine blades. Calculate the maximum theoretical power that can be generated by the turbine.

Gas constant 𝑅 = 8.314 JK-1mol-1

Constant pressure molar heat capacity of nitrogen:𝐶𝑝,𝑚 = 29.125 JK-1mol-1


1
Expert's answer
2021-01-10T18:25:30-0500

"P=\frac At=\frac{mgh}{t}=\frac mt gh=500\cdot 9.8\cdot 7.5=36750 ~W."